OpenStudy (fools101):

Zeros of Polynomial Functions Other one please help! y = 1/4x4 - 3x2

OpenStudy (anonymous):

x = 0 is a zero

OpenStudy (anonymous):

\[{ 1 \over 4 } x^4 -3x^2\] \[x^2({1 \over 4 } x^2 - 3)\] \[x = 0\] OR \[{1 \over 4}x^2 -3 = 0\] \[x^2=12\] \[x = 2\sqrt{3}\]OR \[x=-2\sqrt{3}\]

jimthompson5910 (jim_thompson5910):

y = 1/4x^4 - 3x^2 0 = 1/4x^4 - 3x^2 1/4x^4 - 3x^2 = 0 x^4-12x^2 = 0 z^2 - 12z = 0 ... note: z = x^2 z(z-12) = 0 z = 0 or z-12 =0 z = 0 or z = 12 x^2 = 0 or x^2 = 12 x = 0, x = sqrt(12) or x = -sqrt(12) x = 0, x = 2*sqrt(3) or x = -2*sqrt(3)

OpenStudy (fools101):

Thank u!!! So much!